OpenStudy (anonymous):

Calculate the time required to send a 56 byte packet from Sydney to Melbourne if the data rate is155mbps, the distance traveled is 1000km and the speed of propagation is 2.8*10^8m/s.

OpenStudy (barrycarter):

Start by figuring out how long it takes the first byte to get there.

OpenStudy (anonymous):

56*8=448bits 448/(155*10^6)=2.89*10^-6 (2.8*10^8m/s) / (1*10^6)= 280s 280s*2.8*10^8=8.092*10^-4

OpenStudy (anonymous):

is this right????

OpenStudy (phi):

 I am guessing the length of time of the packet is 2.88 microsec, but it takes about 3.57 millisec to travel 1000 km. so add them together (but 3 microsec doesn't really affect things very much)